[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Tue Jul 11, 2017

HSU Shift Log nhc.tsbadmin at noaa.gov
Tue Jul 11 08:59:50 UTC 2017


HSU Shift Log for Tue Jul 11 2017 04:59:44 GMT-0400 (EDT)

Shift: A

Forecaster: R Berg

Technical Issues Summary:

ATCF issues:

N-AWIPS issues: ASCAT stopped coming in last evening. Called the SDM, who  
determined that it was a problem on NESDIS's end. Data began flowing again  
during the midnight shift.

AWIPS-2/NCP issues:

Other Techincal issues:

Miscellaneous issues:
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170711/5035440d/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list